### v2

# Official-NV: An LLM-Generated News Video Dataset for Multimodal Fake News Detection

Abstract
News media, especially video news media, have penetrated into every aspect of daily life, which also brings the risk of fake news. Therefore, multimodal fake news detection has recently garnered increased attention. However, the existing datasets are comprised of user-uploaded videos and contain an excess amounts of superfluous data, which introduces noise into the model training process. To address this issue, we construct a dataset named Official-NV, comprising officially published news videos. The crawl officially published videos are augmented through the use of LLMs-based generation and manual verification, thereby expanding the dataset. Furthermore, the proposed dataset is benchmarked against several baselines to demonstrate its effectiveness in multimodal news detection.
Index Terms—  Multimodal Fake News Detection, News Video Dataset, Large Language Model, Data Augmentation
1 INTRODUCTION
With the video-oriented trend of news media and the rise of video platforms, multimodal fake news has become increasingly prevalent. This type of fake news mixes text, images, and videos to mislead the public [1, 2, 3, 4, 5]. Detecting fake news videos manually is labor-intensive and inefficient, which is insufficient for the large number of emerging video platforms. Therefore, Multimodal Fake News Detection (MFND) [6, 2, 7, 8] is a significant subtask of Natural Language Processing (NLP) and Computer Vision (CV). By analyzing multiple modalities, we can identify and flag fabricated news more accurately, enhancing the reliability of online content.
Fig. 1: Three examples of fake news videos. From top to bottom, the title is inconsistent with the other two modalities,the speech text is inconsistent with the other two modalities, and the video frames are inconsistent with the other two modalities.
Numerous datasets have been constructed to address fake news detection. Some existing MFND datasets include FVC-2018[9], FakeNewsNet[10], BuzzFace[11], VAVD[12], Liar[13], Weibo[5, 14, 15], and MCFEND[8]. These datasets contain several modalities of information, such as images, text, metadata, and social context. In recent years, the rise of video social media has shifted the focus of MFND to the video modality, so researchers have proposed some MFND datasets for videos. Liu et al.[16] presented a dataset called COVID-VTS, which contains 10k English videos posted by verified users from Twitter, and proposed a fine-grained fact-checking system. Qi et al.[3] collected a total of 5538 Chinese short videos posted by users under 738 events from two short video platforms, Douyin and Kuaishou, and established a video dataset named FakeSV. It consists of 1827 real videos, 1827 fake videos, and 1884 debunked videos. Each record includes videos, titles, metadata, comments, and user information.
However, these fake news video datasets consist of unofficial, user-uploaded videos that often contain numerous duplicates or video memes. Our objective is to train on a dataset with reduced noise and enhanced quality, thereby enabling the model to achieve superior performance. To tackle this challenge, we introduce Official News Videos (Official-NV for short), a dataset comprising officially published English news videos. Specifically, Official-NV encompasses a total of 10,000 videos, evenly distributed between 5,000 authentic and 5,000 fabricated videos. Each video incorporates at least two modalities: the title and video frames. Additionally, videos featuring voice-overs also include the speech text modality. We classify a video as a fake news video when there is a discrepancy among its modalities. Fig. 1 illustrates three examples of fake news videos.
The main contributions are summarized as follows:
  • $\bullet$
    We present Official-NV, a dataset consisting of officially published English news videos. It has less noise and higher video quality compared to other datasets.
  • $\bullet$
    We leverage a Large Language Model (LLM) to generate texts by synonym or antonym substitution, and expand the dataset efficiently.
  • $\bullet$
    We conduct grouped experiments and ablation studies on Official-NV. The experimental results demonstrate that several baseline models have good performance on the proposed dataset.
Table 1: Examples of texts that are before and after modification
Original Text Modificated Text
China’s booming tea industry imbued with new momentum China’s tea industry surges forward with rejuvenated vitality (TT)
China seeks to build world’s largest national park system China aims to dismantle extensive national park network (FT)
…we have done few the major infrastructure projects with the under BRI initiative… … a few significant infrastructure projects have been accomplished under the BRI initiative… (TS)
…human rights experts and scholars from China and other countries exchanged their views on the contributions of the BRI cooperation in terms of global development and human rights progress… …human rights experts and scholars from China and other nations expressed their concerns over the limitations of the BRI cooperation in terms of global development and human rights progress… (FS)
2 THE OFFICIAL-NV DATASET
In this section, we provide a detailed introduction to the construction process of Official-NV. Furthermore, we conduct data analysis and compare the differences between Official-NV and other MFND datasets.
2.1 DATA COLLECTION
Videos in most other MFND datasets are sourced from user-generated content platforms, such as TikTok, Twitter, and YouTube. Consequently, they often include numerous duplicate remakes or memes lacking analytical value, introducing noise into our model training. To address this, we select Xinhua as our target platform for acquiring videos. Xinhua publishes content from official media and is a comprehensive news information service portal hosted by the Xinhua News Agency. It stands as the most influential online media in China and operates as a multi-language website with a global audience. Specifically, we utilize web crawler technology to gather news videos spanning from 2021 to 2024, focusing on Xinhua’s three primary categories: “China”, “The World”, and “Others”. To minimize noise, we manually screen and remove all videos shorter than 15 seconds. Following this, we compute the cosine similarity between every pair of videos in the same category [17], eliminating those with a similarity score exceeding 0.8. This process result in a curated collection of 2,500 news videos, classified as “True Origin” (TO). Within this dataset, 1,000 news videos feature three modalities: title, speech text, and video frame, the other 1,500 videos containing only two modalities, namely the title and video frame, due to the absence of dubbing. The information presented across these three modalities in the TO category remains consistent.
Table 2: Example of prompt and LLM responses
Type Content
Prompt Title: “Turkiye to introduce anti-dumping measures on solar panel imports from 5 countries”, choose a suitable Angle from the direction, position, quantity, action, object and time change. Change the title to another expression with the opposite meaning.
ChatGPT4o Title: “Turkiye Lifts Trade Restrictions on Solar Panel Imports from 5 Countries” This new title presents the opposite meaning by focusing on lifting restrictions rather than imposing anti-dumping measures.
2.2 Data Augmentation
Inspired by methods for generating inconsistent texts and LLM image editing to extend datasets[16, 4], we leverage LLMs to generate fake news data and expand upon real news data after collecting 2,500 news videos. We perform data augmentation for text and video frames using two approaches:
Title and Speech Text. After comparing the generation effects of ERNIE, Qwen, Llama, GPT, etc., we choose to use ChatGPT4o to generate new data. Specifically, we provide prompts to ChatGPT, asking it to alter the given text to convey the same or a completely different meaning by changing aspects such as direction, position, quantity, action, object, and time. An example of such a conversation is presented in Table 2. Following this, we manually remove any redundant information from the responses and screened out poorly generated instances. The altered titles, either retaining the original meaning or conveying a new one, are classified as “True Title” (TT) and “Fake Title” (FT), respectively. Similarly, modified speech texts are categorized as “True Speech” (TS) and “Fake Speech” (FS). Table 1 displays some examples of the data after LLM generation and manual adjustment.
Table 3: Comparison of datasets for MFND
Name Language Source Instances (Fake/Real) Title (Avg.length) Speech (Avg.length) Video (Avg.length) Publisher
COVID-VTS English Twitter 5000/5000 19.2 69.2 26.5s Verified Users
FakeSV Chinese Douyin,Kuaishou 1827/1827 32.7 167.4 38.5s Users
Official-NV English Xinhua 5000/5000 11.3 126.6 69.2s Official Media
Fig. 2: Distributions of news videos
Video Frame. We adopt the same method used in COVID-VTS to create inconsistent videos. We compute the cosine similarity between videos to identify the one most similar to the target video and then replace all of the target video’s frames with frames from this similar video, while keeping its original title and speech text intact. These videos are labeled as “Fake Frame” (FF).
Following data augmentation for each video, we end up with six categories of videos. We prioritize preserving the content generated by LLM and select two instances from the True categories and two from the Fake categories from among these six, resulting in a final dataset comprising a total of 10,000 entries.
2.3 Data Analysis
The Official-NV dataset comprises 10,000 videos, derived from 2,500 official Xinhua news videos through LLM data augmentation. Among these, 4,000 videos include the title, speech text, and video frames, while 6,000 videos feature only the title and video frames due to the absence of dubbing. The dataset maintains an equal number of positive and negative samples, with 5,000 each. Depending on the data generation method, the content can be segmented into six distinct categories. Alternatively, based on Xinhua’s news classification during video acquisition, the data is divisible into three categories: “China”, “The World”, and “Others”. The distribution of these videos is illustrated in Fig. 2.
When compared to other MFND video datasets, Official-NV offers notable advantages. In our LLM text generation process for data augmentation, we utilize a word replacement strategy, swapping words with similar or contrasting meanings. This ensures that the length of both positive and negative texts remains comparable, thereby preventing significant length disparities. With an average video duration of 69.2 seconds, Official-NV boasts considerably longer videos than other datasets, allowing for a richer information content per video. Unlike datasets sourced from user-generated platforms, Official-NV is curated from officially published Xinhua content, yielding videos of superior quality and reduced noise. A comparative analysis of MFND datasets is presented in Table 3.
In this section, we conduct experiments to evaluate the performance of representative fake news detection methods on our newly proposed Official-NV dataset. Specifically, we aim to answer the following Evaluation Questions (EQs):
EQ1: Can the existing method demonstrate its effectiveness on the Official-NV dataset and maintain its performance across several different data augmentation techniques?
EQ2: What is the significance of different modalities in fake news detection?
2.4 Experimental Settings
To address the aforementioned evaluation challenges, we devise the following experimental settings:
Prior to the experiment, we extract key features from the title and voice text of each video in the dataset using BERT[18] and BART[19], respectively. Additionally, we employ CLIP[20] to encode the visual features present in the video frames, thereby enabling a comprehensive representation of the visual content.
During the experimental phase, we select AdamW[21] as our optimizer, cross-entropy as our loss function, and accuracy, F1 score, precision, and recall as our evaluation metrics. This approach allow us to obtain a thorough understanding of the dataset’s performance. Furthermore, we implement 5-fold cross-validation to enhance the reliability and reproducibility of our findings.
2.5 Baselines
To establish a robust and comprehensive benchmark, we conduct experiments using three representative baseline models: BART[19] is a pre-training model designed for denoising sequence-to-sequence tasks, combining the strengths of both bidirectional and autoregressive Transformers. SV-FEND[3] employs cross-modal correlation to select the most informative features and incorporates social context information for detecting false news. TwtrDetective[16] integrates cross-media consistency checking to identify token-level malicious tampering across different modalities and provides explanations for its detections.
Table 4: Results of grouped experiments
Methods Data Acc. F1 Pre. Recall
BART Group T Group S Group F All 87.78 78.79 62.99 78.20 87.71 78.47 62.42 78.19 87.76 78.32 62.47 78.25 87.66 78.92 62.87 78.20
SV-FEND Group T Group S Group F All 85.42 76.06 62.04 78.00 85.35 74.53 60.57 78.00 85.35 76.19 75.87 78.02 85.35 74.03 68.00 78.00
TwtrDetective Group T Group S Group F All 86.28 77.88 64.22 78.70 86.19 75.88 61.99 78.70 86.26 79.52 62.56 78.70 86.14 75.18 61.85 78.70
2.6 Grouped Evaluation (EQ1)
We utilize three baseline models to address EQ1 by conducting experiments on the Official-NV dataset. Specifically, we divide the dataset into three groups: Group T, consisting of videos with modified titles (i.e., all videos labeled as TT and FT); Group S, comprising videos with modified speech text (i.e., all videos labeled as TS and FS); and Group F, which includes videos with modified video frames and all real videos crawled from Xinhua (i.e., all videos labeled as TO and FF). We perform experiments using the three baseline models on these three groups of data as well as on the entire dataset, evaluating the performance of the models across different modification methods and the overall dataset. Our results are present in Table 4, from which we can draw the following conclusions:
1) The accuracy of the three baseline models is above 0.78 on all Official-NV dataset, which is higher than the accuracy achieved on COVID-VTS in which the best accuracy is 0.71, demonstrating that the commonly used MFND model can have good effect on Official-NV dataset. It also serves as additional evidence that the videos in Official-NV exhibit superior quality and contain less noise;
2) Among the three groups, Group T yields the most impressive experimental results, while Group F performs the least favorably. This disparity suggests that the models are more responsive to alterations in short text and less sensitive to changes in imagery compared to textual modifications;
3) Despite Group F’s lower performance, its accuracy remains above 0.62. This observation indicates that the baseline models can sustain their performance even when subjected to different data augmentation techniques.
Table 5: Results of ablation experiments
Methods Acc. F1 Pre. Recall
BART 78.20 78.19 78.25 78.20
w/o T w/o S w/o F w/o T&S w/o T&F w/o S&F 58.80 69.40 75.80 50.00 57.80 69.00 58.38 68.64 75.62 33.33 50.77 67.30 59.17 71.49 76.59 25.00 68.20 73.99 58.80 69.40 75.80 50.00 57.80 69.00
SV-FEND 78.00 78.00 78.02 78.00
w/o T w/o S w/o F w/o T&S w/o T&F w/o S&F 60.10 70.30 76.20 53.20 57.50 68.60 58.65 69.93 76.20 53.18 55.40 68.14 61.74 71.35 76.21 53.21 59.24 69.74 60.10 70.30 76.20 53.20 57.50 68.60
TwtrDetective 78.70 78.70 78.70 78.70
w/o T w/o S w/o F w/o T&S w/o T&F w/o S&F 60.10 70.90 76.30 53.30 57.90 68.70 57.97 70.29 76.27 52.12 56.02 67.03 62.66 72.76 76.44 53.66 59.53 73.45 60.10 70.90 76.30 53.30 57.90 68.70
2.7 Ablation Study (EQ2)
To address EQ2, we conduct ablation studies. Specifically, we perform experiments on the T, S, and F data by removing one and two modalities from the dataset, respectively. Since SV-FEND and TwtrDetective utilize cross-modal attention for multi-modal information fusion, the corresponding cross-modal attention layers are removed when a modality was absent. The results of these experiments are presented in Table 5.
From the results of the ablation studies, we can know:
1) Modality ablation has some impact on the performance of the three baseline models. The influence of different modalities on model performance can be ranked as Title >Speech Text >Video Frame;
2) None of the three baseline models performs well when dealing with the single video frame modality. This is because negative samples are created by replacing frames from other videos without altering the original frames, highlighting the unique challenges presented by this dataset.
3 CONCLUSION
We propose a video dataset for MFND, constructed using LLM data augmentation. Following this, we furnish a comprehensive description of the creation process of Official-NV and delve into the specifics of the data entities’ attributes and distribution. To assess the dataset’s efficacy, we carry out numerical experiments, evaluating the performance of various baselines and exploring the significance of each modality. Upon comparing the results, it becomes evident that the baselines exhibit strong performance on Official-NV, with the text modality proving to be particularly influential. Looking ahead, we intend to explore additional avenues that hold promise in enhancing the predictive power of this dataset, with the aspiration that it will pave the way for further advancements in MFND research.